# SingingHead: A Large-scale 4D Dataset for Singing Head Animation

### v2

@@ 5.2 Compare with 3D Facial Animation Methods
Figure 6: Qualitative comparison with 2D talking head methods. Our method can generate singing videos with higher quality and accuracy in terms of facial movements especially in the mouth region, identity preservation, and video quality.
Furthermore, previous methods model the audio-driven 3D facial animation as a deterministic regression task, which can only generate one results given a input audio. In contrast, our method models it as a generation task, which can generate diverse facial motions according to one input audio. We generate $n$ samples of result for each testing audio by sampling the latent gaussian distribution of our model for $n$ times, and report the evaluation results under 1, 10, 30 samples. The outstanding performance of our method on the average pairwise distance (APD) illustrates the diversity of 3D facial motion generated by our method.
5.3 Compare with 2D Talking Head Methods
In addition to 3D facial motion generation, our framework UniSinger can also render the generated 3D facial motion to 2D singing video given an additional reference image. The whole process is to take audio and a reference image as input and synthesize a singing portrait video, which consists of the 2D talking head task. Therefore, we evaluate our method by comparing it to several SOTA 2D talking head methods [60, 51, 58].
To prove the necessity of a singing-specific dataset in the task of singing head animation, we need to show that talking head methods trained on a large-scale talking head dataset can not generalize well to the singing scenario. For this purpose, we select two typical talking head methods: PC-AVS [60] and SadTalker [58], and use their pre-trained models which are trained on large-scale talking head videos to directly make inferences on our SiningHead test set. We show the qualitative results and quantitative results in Fig. 6 and Tab. 3 respectively. We can see that our method achieves the best performance on accuracy and amplitude of facial movements, especially on the mouth region, which demonstrates the importance and necessity of a singing-specific dataset in singing head animation. Moreover, we choose an additional one-shot talking head generation method Audio2Head [51] to show the superiority of our approach itself. As shown in Fig. 6 and Tab. 3, our method achieves promising performance on the accuracy and quality of the generated portrait singing video.
Method LMD $\downarrow$ FID $\downarrow$ SSIM $\uparrow$
PC-AVS [60] 7.372 114.7 0.580
Audio2Head [51] 5.353 51.13 0.681
SadTalker [58] 5.895 19.54 0.727
Ours 3.996 13.33 0.717
Table 3: Comparison with 2D talking head methods. We report the landmark distance of the mouth region, FID, and SSIM of the generated singing videos of the selected 10 identities.
6 Conclusion
@@
In this paper, we focus on the task of singing head animation, and collect the first large-scale high-quality 4D facial animation dataset specialized for singing, named SingingHead. This dataset contains more than 27 hours’ synchronized portrait video, 3D facial motion sequence, singing audio, and background music from 76 subjects. Along with the dataset, we propose a unified framework called UniSinger to solve both the 3D and 2D singing head animation tasks. Compared with SOTA methods, our framework achieves superior performance on both 3D singing facial motion generation and 2D singing portrait video synthesizing.

### v3

@@ V-A Implementation details
Figure 10: Qualitative results of 2D singing portrait video synthesis. Our method can generate singing portrait videos with higher quality and accuracy in terms of facial movements, especially in the mouth region, identity preservation, and video quality.
Figure 11: Qualitative results of 3D singing head animation. We show the generated 3D facial motion sequences according to two input singing audios. Our method can generate more accurate and natural motion sequences compared to other methods. The pre-trained models perform the worst, which demonstrates the necessity of our singing-specific dataset for the 3D singing head animation task.
V-C 2D Singing Portrait Video Synthesis
Evaluation metrics. We evaluate the performance of 2D singing portrait video synthesis from three aspects. (1) Video quality. We use full-reference metrics, including Frechet Inception Distance (FID) [89, 90], Learned Perceptual Image Patch Similarity (LPIPS) [91] and Structural Similarity (SSIM) [92], to measure the realism, perceptual and structural similarity of the generated video frames compared to the ground truth. Besides, the no-reference metric Cumulative Probability Blur Detection (CPBD) [93] is adopted to evaluate the sharpness of the generated frames. (2) Accuracy of facial movements. We adopt Landmark Distance (LMD) [94] to evaluate the alignment of generated facial movements with the ground truth, thereby assessing the synchronization of the generated portrait video with the input audio. Specifically, we first detect the 68 facial landmarks in both generated and ground truth frames using [95]. Then, we calculate the distance between the generated and ground truth landmarks in the mouth region and the entire face region to obtain the metrics M-LMD and F-LMD, respectively. (3) Alignment with audio. To measure whether the facial movements in the generated video align well with the input audio, we adopt the Beat Align Score (BA) [88] that is applied to the 68 facial landmarks.
Baseline methods. We select some classic 2D talking head methods including MakeItTalk [25], Audio2Head [15], SadTalker [1], EAT [16] and DreamTalk [17]. Since the 2D talking head methods are typically trained on large-scale datasets and have zero-shot capability, we directly evaluate their pre-trained models on the selected identities of the SingingHead test set.
The quantitative and qualitative results are shown in Tab. IV and Fig. 10, respectively. We can see that our method achieves much more accurate facial movements and better alignment with input audio. Although other pre-trained models are trained on very large-scale talking head datasets, they often fail to generalize well to the singing task. As a result, these methods struggle with lip synchronization, especially in scenarios where the mouth needs to be opened wider, which is not common in regular talking. Therefore, it is essential to use singing-specific data to help existing 2D talking head methods generalize well to the singing task and achieve better performance in 2D singing portrait video synthesis. Moreover, our method also achieves higher video quality and more realistic and vivid visualization results.
VI Conclusion
@@
In this paper, we focus on the task of singing head animation and collect the first large-scale high-quality 4D facial animation dataset specialized for singing, named SingingHead. This dataset contains more than 27 hours’ synchronized portrait video, 3D facial motion sequence, singing audio, and background music from 76 subjects. Along with the dataset, we benchmark existing audio-driven 3D facial animation methods and 2D talking head methods on the singing task. Moreover, we propose a unified framework named UniSinger to achieve both singing audio-driven 3D singing head animation and 2D singing portrait video synthesis, which achieves competitive results on both 3D and 2D benchmarks.